OpenStudy (anonymous):

Use a related equation that has the variable on one side. Then simplify the other side. d/6=2 A.d=12 B.d=8 C.d=4 D. d=1/3

OpenStudy (anonymous):

i think its d

OpenStudy (anonymous):

because 3/1 = 3 and 3 goes in 6 2 times

OpenStudy (anonymous):

ans is A

OpenStudy (anonymous):

d/6=2 d=2*6=12

OpenStudy (anonymous):

thx